7. A ball is thrown vertically upward with a speed of 7.1 m/s. Find:


(9 points)


a. Maximum height reached the ball


b. Its time of flight


c. Velocity the ball would return to its starting point

1
Expert's answer
2022-05-02T09:33:02-0400

a.


"h_{max}=\dfrac{v_0^2}{2g}=\dfrac{(7.1m/s)^2}{9.8m/s^2}\approx2.572m"

b.


"h(t)=v_0t-\dfrac{gt^2}{2}""h=0=>v_0t-\dfrac{gt^2}{2}=0"

"t_1=0, t_2=\dfrac{2v_0}{g}"

"t_2=\dfrac{2(7.1m/s)}{9.8m/s^2}\approx1.45s"

c.


"v_{ground}=-v_0=-7.1m\/s"